1. In 8:00 AM, the population of a bacteria is 1000. At 10:30 AM, the number of bacteria triples. At what time will the population become 100 times the initial population of bacteria? What will be the population at 2:00 PM?  Ans. 6:29 PM; 13,967
  2.  If the present population of a certain country is 40 million and in 10 years, the population is 50 million, what will be its population 20 years from now? Ans. 62.5 million
1
Expert's answer
2021-11-14T16:22:48-0500

1) Bacteria at initial state=1000,8:00am

Bacteria at final state=100000

Tripled after 2.5hrs

The exponential form is;

"100000=1000*3^{(\frac{t}{150})}"

"100=3^{(\frac{t}{150})}"

"ln\ 100=\frac{t}{150}ln\ 3"

"\frac{t}{150}=(\frac{ln\ 100}{ln\ 3})"

"{t}=(\frac{ln\ 100}{ln\ 3})*150"

"t=628.77mins \approx629mins"

"t=10hrs \ 29 mins"

Time =8.00 am+10hr 29 mins

=6.29pm


Population at 2.00pm

"P=1000*3^{(\frac{t}{150})}"

t=2.00pm -8.00am=6hrs=360mins

"P=1000*3^{(\frac{360}{150})}"

=41899.83


2)"\frac{dN}{dt}=kn"

"\int \frac{dN}{N}=\int kdt"

"ln\ N=kt+C"

Given for t=0, N=40million

"\implies ln(40)=C"

And for t=10, N=50million

ln(50)=k*10+ln(40)

"k=\frac{1}{10}(ln\ 50-ln\ 40)=\frac{1}{10}ln\ (1.25)"

Hence, "ln\ (N)=\frac{1}{10}ln\ (1.25)t+ln\ (40)"

For t=20

"ln\ (N)=\frac{1}{10}ln\ (1.25)*20+ln\ (40)"

="ln\ (1.25)^2+ln\ (40)=ln\ (40*1.25^2)"

N=40*1.252

=62.5million